OpenStudy (anonymous):

ODE second order d^2 y/dx^2 = y'' = xy find the general solution, please help, i dont know how to do.. i have already forgotten the method.. need solution

OpenStudy (anonymous):

I wonder if it is separable? \[ \begin{array}{rcl} \frac{d^2y}{dx^2} &=& xy \\ \frac{dy}{dx}\frac{dy}{dx} &=& xy \\ \frac{1}{y}\frac{dy}{dx}\frac{dy}{dx} &=& x \\ \int \int \frac{1}{y}\frac{dy}{dx}\frac{dy}{dx}dxdx &=& \int \int x dx dx\\ \end{array} \]

OpenStudy (anonymous):

Since \[ dy = \frac{dy}{dx}dx \]I'm thinking \[ \frac{dy}{dx}\frac{dy}{dx}dxdx = \frac{dy}{dx}dydx = \frac{dy}{dx}dxdy = dydy \]

OpenStudy (anonymous):

If so then don't we just have: \[ \int \int \frac{1}{y}dydy = \int \int xdxdx \]

OpenStudy (anonymous):

@ikamashi Can you do these integrals?

OpenStudy (abb0t):

What you have is: y'' - xy = 0 since: \[\frac{ d^2y }{ dx^2 } = y''\]

OpenStudy (abb0t):

You can also use series to solve this. Eulers method.

OpenStudy (abb0t):

I am going to assume what I stated above, that y''-xy = 0 and use series Using: \[y(x) = \sum_{n=0}^{∞}a_nx^n\] Then I am going to find y' and y'' I am sure that you are able to do these derivatives easily so I am going to leave it to you to figure out the derivatives [NOTE: a and n are both constants] Next, you are going to plug in y and y'' into the differential [NOTE: make sure to shift down the 2nd differential, you want it as n=0 instead of n=2 to get the series in terms of x^n]

OpenStudy (anonymous):

is \[\frac{ d ^{2}y }{ dx ^{2} }=\frac{ dy }{ dx }\frac{ dy }{ dx }\] ??? i got \[y (\ln y-1) =\frac{ x ^{3} }{ 6 }+ c _{i}(x-y) \]

OpenStudy (anonymous):

anyway, thx wio n abb0t, for helping me

OpenStudy (experimentx):

you should be using power series, the solution is in one of the standard special function http://en.wikipedia.org/wiki/Airy_function